40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Wed Aug 2 03:35:18 2017, Zach, Electronics, SiFi, Current driver range increased to 200 mA SiFi_current_driver2_coarse_response_8_1_17.pngSiFi_current_driver2_fine_response_8_1_17.pngR31.jpg
    Reply  Wed Aug 2 08:31:58 2017, Johannes, Electronics, SiFi, Current driver range increased to 200 mA 
       Reply  Wed Aug 2 17:23:48 2017, Zach, Electronics, SiFi, Current driver range increased to 200 mA SN104978_PvsI.pdf
    Reply  Thu Aug 3 18:06:05 2017, Zach, Update, SiFi, More system characterization measurements (laser drivers, heads) SiFi_laser_PvsI_8_2_17.png
Message ID: 1654     Entry time: Wed Aug 2 17:23:48 2017     In reply to: 1653
Author: Zach 
Type: Electronics 
Category: SiFi 
Subject: Current driver range increased to 200 mA 

(We discussed this earlier today, but I wanted to reply on here for others...)

My RIO lasers are actually specified to be run at higher current, it seems. One unit has Iop = 150 mA / Imax = 160 mA; the other Iop = 140 mA / Imax = 155 mA.

As shown below, one of the lasers barely even turns on below 100 mA:

The full laser head datasheets and all characterization data are/will be available on the new SiFi wiki page: https://nodus.ligo.caltech.edu:30889/CryoWiki/doku.php?id=projects:sifi

I'm planning to implement the stacked-diode voltage limiter that you reference, as well. I looked into getting TO-220 power resistors with R = 68 ohm, so that the output current wouldn't exceed the RIO max spec, but these don't appear to be easy to find.

Quote:

I'm not sure if it's a good idea to run the laser diodes at a significantly higher current than 100mA. Their recommended factory setting is somewhere around 100 mA depending on the serial number, and the maximum recommended bias current is usually ~110%-115% of that value. We have received three current drivers from Rich that can go up to 200mA, so there's also a spare one on top of one of the racks. We also had two laser units fail with those drivers, and the exact failure mode was never completely narrowed down, but temporarily running too much current could definitely have played a role. We also made a voltage limiting modification to the coating noise cavity laser drivers using a series of four diodes in parallel to the laser diode that turn active when the drive voltage becomes higher than 5-6V that you can see in this post.

 

Attachment 1: SN104978_PvsI.pdf  386 kB  | Hide | Hide all
SN104978_PvsI.pdf
ELOG V3.1.3-